Joan 3 times as old as her sister. In 3 years, she will be two more than twice the age of her sister will be then. What are their ages? Using variable g, express your answer to the following. (use small letter and no space/s in between)
a. Sister's age
b. Joan's age
c. Sister's age in three years
d. Joan ages in three years
e. Equation
ANSWER:
a. g
b. 3g
c. g+3
d. 3g+3
however in letter e (equation) how do i write it correctly?
(3g+3) = 2+2(g+3) or 3g+3 = 2+2(g+3) or 2+2(g+3) ?

Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52776)

.

a.  g    (<<<---=== the sister's age)


b.  3g


c.  g+3


d.  3g+3


e.  equation  3g + 3 = 2 + 2*(g+3)

              3g + 3 = 2 + 2g + 6

              3g - 2g = 2 + 6 - 3

                 g    = 5.


ANSWER.  The sister is 5 years old.  Joan is 15 years old.


CHECK.   In 3 years, Joan will be 15+3 = 18 years old;  the sister will be 8 years old.

         18 = 2 + 2*8.        ! Correct !

Solved.

Is everything clear to you in my solution ?

---------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.



Answer by josgarithmetic(39616)

duplicate - solved two days ago